OpenStudy (anonymous):

Can please help me 7/10m - 17 = -1/2m -5

OpenStudy (jennychan12):

7/10m - 17 = -1/2m -5 5/6m = 12 5m = 72 m = 72/5 = 14.4

OpenStudy (jennychan12):

i'm a little worried that you don't know your algebra... or do you just hate fractions (it's ok, i do too!)

OpenStudy (jennychan12):

oh sorry. it's just you'll need to brush up on your algebra. cuz this is just solving stuff. it'll be harder as you move on... just saying...

OpenStudy (jennychan12):

wait what???? i never learned this in prealgebra.... i learned this is algebra ab.. but still...

hero (hero):

Wow, this conversation reminds me of comic view

OpenStudy (anonymous):

@Hero what is comic view ..............??????????

hero (hero):

It's a show where comedians say funny stuff and people laugh at them.

OpenStudy (anonymous):

ok but can you help me

hero (hero):

I already helped you which apparently did not help you since you still do not know how to add fractions.

OpenStudy (anonymous):